OpenStudy (anonymous):

find how many ways the integers 1,2,3,4,5,6,7,8 can be placed in a circle if: - Atleast three odd numbers are together?

OpenStudy (anonymous):

what do you mean, placed in a circle?? what type of q is this?? trigo, geometry, algebra, what??

OpenStudy (anonymous):

probability

OpenStudy (anonymous):

This from the internet Q2. Either the 4 odd numbers are together or 3 are together, and the 4th is separated from them. In the first case, the 4 odd numbers can be arranged in an odd block in P(4,4) = 24 ways. The remaining 4 even numbers can then be arranged in an even block in P(4,4) = 4! ways also. In the second case, 3 of the 4 odd numbers can be arranged in the odd block in P(4,3) = 4!/(4-3)! = 24 ways. The 4 even numbers can be arranged in an even block in P(4,4) = 24 ways. The 4th odd number can then be wedged into one of the 3 spaces between adjacent even numbers in C(3,1) = 3 ways. So the number of ways to arrange at least 3 odd numbers together is P(4,4)·P(4,4) + P(4,3)·P(4,4)·C(3,1) = 24·24 + 24·24·3 = 2,304.

OpenStudy (anonymous):

well

OpenStudy (anonymous):

thats that? lolol

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

omg i hate this question lol

OpenStudy (anonymous):

I'm not saying it's right, u better check it, I hate numbers...

OpenStudy (anonymous):

no the answer is right i just don't get it

OpenStudy (anonymous):

what does it mean the 4th is separated?

OpenStudy (anonymous):

Question says "at least 3" so u have to consider the cases when 3 are together and 4 are together.

OpenStudy (anonymous):

i understand that part what i did was just (4!*4!) + (3!*5!) obviously it was wrong lol

OpenStudy (anonymous):

This sort of question is more like a puzzle really, the counting is just math, the reasoning is something else. Best to read it over a few times, let it sink in, why it works...

OpenStudy (anonymous):

kk ill do that thanks:)

OpenStudy (anonymous):

ur welcome.